# Timer-XL: Long-Context Transformers for Unified Time Series Forecasting

### v3

@@ Abstract
We present Timer-XL, a generative Transformer for unified time series forecasting. To uniformly predict 1D and 2D time series, we generalize next token prediction, predominantly adopted for causal generation of 1D sequences, to multivariate next token prediction. The proposed paradigm uniformly formulates various forecasting scenarios as a long-context generation problem. We opt for the generative Transformer, which can capture global-range and causal dependencies while providing contextual flexibility, to implement unified forecasting on univariate series characterized by non-stationarity, multivariate time series with complicated dynamics and correlations, and covariate-informed contexts that include both endogenous and exogenous time series. Technically, we propose a universal TimeAttention to facilitate generative Transformers on multiple time series, which can effectively capture fine-grained intra- and inter-series dependencies of flattened time series tokens (patches), and is further enhanced by deftly designed position embeddings for the temporal and variable dimensions. Timer-XL achieves state-of-the-art performance across challenging forecasting benchmarks through a unified approach. Based on large-scale pre-training, Timer-XL also demonstrates notable zero-shot performance, making it a promising architecture for large time series models.
@@ 1 Introduction
Transformers have contributed significantly to the fields of natural language and computer vision (Radford et al., 2018; Dosovitskiy et al., 2020), and been extensively applied in time series forecasting, becoming the foundation of specialized forecasters (Zhou et al., 2021; Wu et al., 2021) and large models (Das et al., 2023). As a typical generative task, the quality of predictions heavily relies on the context (Dai et al., 2019). Reliable predictions are made by thoroughly considering endogenous temporal variations and retrieving relevant exogenous correlations into the context (Box, 2013). Further, the pre-training context length, which serves as an indicator of scaling (Kaplan et al., 2020), determines the maximum input and output of generative Transformers, ultimately enabling long-sequence, high-resolution, and high-frequency generation (Yin et al., 2023; Wang et al., 2024a).
@@
However, existing Transformers in the time series field crucially encounter the context bottleneck. As shown in Figure 1, unlike Transformers for natural language and vision that learn dependencies among thousands to millions of tokens (Kirillov et al., 2023; OpenAI, 2023), time-series Transformers typically work around limited contexts of up to hundreds of time series tokens (patches) (Nie et al., 2022). For univariate time series, a short context length leads to an insufficient perception of global tendencies, overlooking widespread non-stationarity in real-world time series (Hyndman, 2018). The excessive reliance on stationarization, such as normalization (Kim et al., 2021), restricts the model capacity and leads to overfitting of Transformers (Liu et al., 2022b). Moreover, instead of regarding multivariate time series as independent channels (Nie et al., 2022), increasing Transformers explicitly capture intra- and inter-channel dependencies (Zhang & Yan, 2022; Liu et al., 2023; 2024a), leading to an urgency of extending the context length to encompass inter-correlated variables.
@@
Recently, generative Transformers, which present a predominant scalable choice of large language models (Zhao et al., 2023) characterized by the decoder-only architecture, have gained increasing attention in the development of large time series models (Rasul et al., 2023; Ansari et al., 2024). Based on their generalization performance and contextual flexibility, one model can accommodate varing input lengths during inference (Liu et al., 2024b). Therefore, pre-training on longer contexts not only empowers them with the fundamental capability to incorporate more contextual information but also enhances the model versatility toward a one-for-all foundation model, which regards any-variate and any-length time series as one context. Even if previous work (Liu et al., 2024a) has achieved unified modeling on flattened tokens in encoder-only Transformers (Nie et al., 2022), our empirical results in Figure 3 reveal that encoder-only Transformers encounter performance bottleneck especially on long-context time series, while generative Transformers can mitigate performance degradation well.
@@
In pursuit of unified time series forecasting, we propose multivariate next token prediction. It unifies forecasting tasks of Figure 1 into the patch-level generation based on long-context sequences. To fully leverage the global-range modeling ability and length-flexibility of generative Transformers, we develop TimeAttention that captures causal temporal dependencies in the channel-dependent approach, presenting a novel masking mechanism from the disentangling perspective. With incorporated relative position embeddings for multivariate series, TimeAttention is aware of the chronological order of time points and achieves permutation-equivariance (Zaheer et al., 2017) of variables. We propose Timer-XL as the extra long version of a generative time-series Transformer (Timer) (Liu et al., 2024c). We enlarge the context to thousands of patch tokens and achieve state-of-the-art in univariate, multivariate, and covariate-informed forecasting. With comparable pre-training scale and model size, Timer-XL outperforms advanced large models in zero-shot forecasting. Our contributions lie in three aspects:
@@
  • We propose multivariate next token prediction and unified time series forecasting, strengthening Transformers with enriched forecasting contexts to make reliable predictions.
@@
  • We introduce TimeAttention, a novel causal self-attention tailored for the time series modality on our proposed paradigm, which enables intra- and inter-series modeling with positional awareness, and maintains the causality and length-flexibility of generative Transformers.
@@
  • We propose Timer-XL, a versatile Transformer for one-for-all forecasting, which mitigates performance degradation in long-context time series, achieves state-of-the-art performance in task-specific benchmarks, and presents notable zero-shot results by pre-training.
@@
Figure 1: We compare the context length (measured in the number of tokens) of Transformers in different modalities and propose Timer-XL that increases the length to thousands of patch tokens. Given the generality across contexts, Timer-XL is a versatile solution for various forecasting tasks.
@@ 2 Related Work
Transformers (Vaswani et al., 2017) for time series forecasting have undergone rapid advancements.
Based on the global-range modeling ability of tokens, Transformers have shown great power in time series, rapidly standing out from RNN-, CNN- and MLP-forecasters, especially on long sequences. Initial Transformer-based forecasters primarily focused on long-term prediction, aiming to extend the forecasting horizon while mitigating the quadratic computational growth associated with increasing sequence length (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021). However, the context (lookback) length of previous models is not growing in pace, which hinders Transformers from making fully informed predictions, producing oversmooth results (Liu et al., 2022b). Meanwhile, another advancement has emerged from univariate to multivariate forecasting. Unlike natural language, time series are high-dimensional and inherently correlated (Hyndman, 2018), further requiring longer contexts that contain relevant endogenous and exogenous variables. To effectively exploit the intra- and inter-series dependencies, tokenization of Transformers has been extensively developed in temporal-wise (Lim et al., 2021), patch-wise (Nie et al., 2022), and variable-wise (Liu et al., 2023) approaches, with deftly designed Transformers for inter-series modeling (Zhang & Yan, 2022; Wang et al., 2024b).
Despite the challenges caused by the limited context for making predictions, few works highlight that these challenges can be uniformly tackled by long-context generative Transformers. Consequently, we leverage the fundamental sequence modeling capability of the vanilla Transformer, extend the context length significantly, and unify various forecasting tasks into multivariate next token prediction.
@@
Recently, time-series Transformers have experienced the evolution from small task-specific models to large foundation models (Das et al., 2023; Woo et al., 2024; Ansari et al., 2024). Among them, decoder-only Transformers, predominantly adopted as the backbone of large language models (Touvron et al., 2023; OpenAI, 2023), have positioned as a scalable and generalizable choice for general time series analysis (Liu et al., 2024c). By independently predicting each token based on preceding contexts, decoder-only models are also multi-length forecasters (Liu et al., 2024b), avoiding resource-intensive training and lookback-search. However, existing decoder-only forecasters are generally trained on a single channel (Nie et al., 2022), making them inaccessible to inter-series dependencies.
@@
Prior work has employed encoder-only Transformers to fully capture dependencies within 2D time series (Liu et al., 2024a). The incompatibility between this architecture and temporal causality may constrain the flexibility and performance of Transformers. To date, the implementation of next token prediction and multivariate time series forecasting in a single Transformer remains a fundamental challenge. Our work addresses this issue by disentangling the fine-grained token dependencies clearly into variable dependencies and temporal causal masks, thereby maintaining temporal causality and capturing inter-series dependencies simultaneously without altering the self-attention mechanism.
@@ 3 Approach
In this section, we first introduce a generative Transformer to illustrate the procedure of next token prediction on 1D time series. As an extension, we design TimeAttention and propose Timer-XL for unified time series forecasting. It is applicable to univariate, multivariate, and covariate-informed scenarios by generalizing the context from 1D sequences to 2D time series.
@@ 3.1 Timer
Timer (Liu et al., 2024c) is a generative Transformer trained by next token prediction (Bengio et al., 2000), which regards single-dimensional time series as non-overlapping patch tokens.
@@
For simplicity, we omit the block index $l$. Timer adopts $\mathbf{W}_{e},\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ that independently embed and project the token embeddings as $\mathbf{H}=\{\mathbf{h}_{i}\}\in\mathbb{R}^{T\times D}$. $\operatorname{TrmBlock}$ includes feed-forward network and self-attention with the temporal causal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$. $\mathbf{h}_{i}\in\mathbb{R}^{D}$ is the contextual representation of the previous $i$ tokens. Predicted $\hat{\mathbf{x}}_{i+1}$ are supervised by ground truth in MSE loss.
@@ 3.2 Generalize 1D Sequences to 2D Time Series
Compared with Equation 2, the multivariate context length increases from $T$ to $NT$. By contrast, the benefit is that this paradigm learns causal dependencies within each sequence while incorporating exogenous variable correlations from other sequences, making it a universal forecasting paradigm that outperforms channel independence or coarse-grained variable-wise modeling experimentally.
Technically, we still adopt the token embedding $\mathbf{W}_{e}\in\mathbb{R}^{D\times P}$ to obtain the patch-wise representation $\mathbf{h}_{m,i}\in\mathbb{R}^{D}$, which will encompass contextual information from $Ni$ tokens through Transformer blocks and be eventually projected by $\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ into the predicted patch token $\hat{\mathbf{x}}_{m,i+1}$.
@@
Position embedding has not been sufficiently explored in time-series Transformers. To avoid the permutation-invariance of self-attention, positional embedding is required to reflect the chronological order of tokens on the temporal dimension. As for the variable dimension, shuffling the input order of variables should not affect anything other than the output order of variables. Formally, the processing on multiple variables should be permutation-equivalent (Zaheer et al., 2017).
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. $\mathbf{R}_{\theta,t}\in\mathbb{R}^{d_{k}\times d_{k}}$ is the rotary matrix with rotation degree $t\cdot\theta$, $\mathds{1}{(}\cdot)$ is the indicator function, and $u,v\in\mathbb{R}$ are learnable parameters for the token to distinguish its endogenous and exogenous variables.
@@
We provide an intuitive example to illustrate the causal dependencies within multivariate time series: considering the 2nd token of time series A. To predict its next token, its representation $\mathbf{h}$ should be exactly dependent on the tokens-$\{1,2,4,5\}$. Similarly, we provide all causal dependencies of each token in Figure 12. Based on the visualized attention mask and variable dependencies presented in Figure 2, where all variables are inter-correlated, we find that causal temporal dependencies in $\mathcal{A}$ can be formally disentangled by the Kronecker product into (1) the adjacency matrix of the variable dependency graph $\mathcal{C}\in\mathbb{R}^{N\times N}$ and (2) the causal temporal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$:
@@
As before, token representations in $\mathbf{H}=\{\mathbf{h}_{m,i}\}\in\mathbb{R}^{NT\times D}$ will be independently processed by feed-forward network and layer normalization, and fed into the next Transformer block.
@@
In multivariate forecasting, the variable dependency forms the complete graph, presenting an all-one matrix $\mathcal{C}$. By generalizing TimeAttention on multiple sequences, generative Transformers can leverage its length-flexibility to encompass relevant covariates as well. In this case, Timer-XL is adapted in two steps: (1) formulate the customized variable dependency as $\mathcal{C}$ and (2) optimize the model using the supervision of target variables. An example (target-$A$-covariate-$B$) of TimeAttention is illustrated on the right of Figure 2. In a nutshell, we adopt position embeddings for the temporal and variable dimensions. To achieve unified time series forecasting, we flatten 2D time series into a unified context and capture fine-grained causal token dependencies.
@@ 4 Experiments
We conduct thorough evaluations of the performance and versatility of Timer-XL. Given that the long-context forecasting paradigm receives less attention in the community, which can be concealed due to the performance saturation on previous benchmarks (Wu et al., 2021; Nie et al., 2022), we delve into the backbone and representation comparison on Transformers. New long-context benchmarks are established, which will be released for the advancement of this field. Detailed datasets, baseline models, and experimental configurations are provided in Appendix B. We also dive into commonly adopted techniques, such as channel independence (Nie et al., 2022) and normalization (Kim et al., 2021). We conclude that generative Transformers can tackle these challenges without specific designs.

### v4

@@ Abstract
We present Timer-XL, a causal Transformer for unified time series forecasting. To uniformly predict multidimensional time series, we generalize next token prediction, predominantly adopted for 1D token sequences, to multivariate next token prediction. The paradigm formulates various forecasting tasks as a long-context prediction problem. We opt for decoder-only Transformers that capture causal dependencies from varying-length contexts for unified forecasting, making predictions on non-stationary univariate time series, multivariate series with complicated dynamics and correlations, as well as covariate-informed contexts that include exogenous variables. Technically, we propose a universal TimeAttention to capture fine-grained intra- and inter-series dependencies of flattened time series tokens (patches), which is further enhanced by deft position embedding for temporal causality and variable equivalence. Timer-XL achieves state-of-the-art performance across task-specific forecasting benchmarks through a unified approach. Based on large-scale pre-training, Timer-XL achieves state-of-the-art zero-shot performance, making it a promising architecture for pre-trained time series models. Code is available at this repository: https://github.com/thuml/Timer-XL.
@@ 1 Introduction
Transformers have been extensively applied to time series forecasting, becoming the backbone of task-specific models (Zhou et al., 2021; Wu et al., 2021) and pre-trained models (Das et al., 2023). While the majority of prior works have focused on long-term forecasting, reliable predictions are made by considering endogenous variations and exogenous correlations in the context (Box, 2013). Besides, the context length of pre-trained Transformers determines the maximum input and output length during inference. Therefore, long-context Transformers are more versatile than shorter ones, facilitating long-sequence and high-resolution generation (Yin et al., 2023; Wang et al., 2024a).
@@
However, existing Transformers in the time series field crucially encounter the context bottleneck. As shown in Figure 1, unlike Transformers for natural language and vision that learn dependencies among thousands to millions of tokens (Kirillov et al., 2023; OpenAI, 2023), time-series Transformers typically operate around limited contexts of up to hundreds of time series tokens (patches) (Nie et al., 2022). For univariate forecasting, a short-context input leads to insufficient learning of global tendencies, struggling to address non-stationarity in real-world time series (Hyndman, 2018). For multivariate forecasting, increasing research has demonstrated the effectiveness of explicitly capturing intra- and inter-channel dependencies (Zhang & Yan, 2022; Liu et al., 2023; 2024a), highlighting the practical urgency of extending the context length to encompass inter-correlated time series.
@@
Recently, causal Transformers characterized by the decoder-only architecture have become a predominant choice of large language models (Zhao et al., 2023) and garnered increasing attention in the development of large time series models (Rasul et al., 2023; Ansari et al., 2024). Based on contextual flexibility and autoregressive next token prediction, one model can accommodate varying lookback and prediction lengths (Liu et al., 2024b). Therefore, pre-training on longer contexts not only empowers them with the fundamental capability to incorporate more contextual information but also enhances the model versatility toward a one-for-all foundation model. Regarding any-variate and any-length time series as one context, previous work (Liu et al., 2024a) has achieved unified modeling on flattened tokens based on noncausal Transformers. However, our empirical results (Figure 3) reveal that encoder-only forecasters may encounter performance degradation in long-context forecasting, while decoder-only Transformers can mitigate this degradation well.
@@
In this work, we generalize the training objective of language modeling to multivariate next token prediction, achieving unified time series forecasting that covers tasks in Figure 1 (right). Based on the decoder-only architecture, we propose TimeAttention to facilitate Transformers on multidimensional time series, presenting Kronecker-based masking mechanism to train time-series Transformers in a channel-dependent approach. With specialized position embedding for multivariate series, TimeAttention is aware of the chronological order of time points and achieves permutation-equivalence (Zaheer et al., 2017) on variables. We enlarge the context to thousands of patch tokens and achieve state-of-the-art on univariate, multivariate, and covariate-informed forecasting benchmarks. By pre-training on large-scale datasets, we present Timer-XL as an extra long version of pre-trained time-series Transformers (Timer) (Liu et al., 2024c), which outperforms recent large models in zero-shot forecasting. Our contributions lie in three aspects:
@@
  • We propose multivariate next token prediction and unified time series forecasting, strengthening Transformers with enlarged contexts to make information-complete predictions.
@@
  • We introduce TimeAttention, a novel causal self-attention tailored for multidimensional time series, facilitating intra- and inter-series modeling with positional awareness and maintaining causality and scalability of Transformers.
@@
  • We propose Timer-XL, a versatile Transformer for one-for-all forecasting, which mitigates performance degradation in long-context time series, achieves state-of-the-art performance in task-specific benchmarks, and presents notable zero-shot performance by pre-training.
@@
Figure 1: We compare the context length (measured by token number) of Transformers in different modalities and propose Timer-XL that increases the length to thousands of patch tokens. Given the generality across contexts, Timer-XL is a versatile solution for various forecasting tasks.
@@ 2 Related Work
Transformers (Vaswani et al., 2017) for time series forecasting have undergone rapid advancements. Initial Transformer-based forecasters primarily focused on long-term forecasting (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021). However, the context (lookback) length is not growing in pace, which hinders Transformers from making information-complete predictions. Another advancement has focused on multivariate forecasting. Unlike natural language, time series are multidimensional and inherently correlated (Hyndman, 2018). To learn intra- and inter-series dependencies, different tokenization of time-series Transformers has been proposed, including point-wise (Lim et al., 2021), patch-wise (Nie et al., 2022), and variable-wise (Liu et al., 2023) approaches, with deftly tailored architectures (Zhang & Yan, 2022; Wang et al., 2024b). However, few works highlight that multidimensional time series can be uniformly tackled by long-context Transformers without architectural modification. In this work, we leverage causal Transformers, which excel at handling long-context sequences, and unify time series forecasting tasks into multivariate next token prediction.
@@
Recently, time-series Transformers have experienced the evolution from small task-specific models to pre-trained large models (Das et al., 2023; Woo et al., 2024; Ansari et al., 2024). Among them, decoder-only Transformer is predominantly adopted as the backbone of large language models (Touvron et al., 2023; OpenAI, 2023), positioning as a scalable choice for general time series analysis (Liu et al., 2024c). By independently predicting each token with supervision, decoder-only models are also multi-length forecasters (Liu et al., 2024b), avoiding resource-intensive training and lookback-search. However, existing decoder-only Transformers are generally pre-trained in a channel-independent approach, making them inaccessible to inter-series dependencies.
@@
Prior work has employed encoder-only Transformers to capture dependencies of multivariate time series (Liu et al., 2024a). However, our empirical study found that this architecture can be incompatible with causal forecasting, limiting the performance of Transformers. To implement next token prediction and multivariate forecasting in a single Transformer, we renovate the attention module, which disentangles fine-grained token dependencies into variable dependencies and temporal causal masks, capturing intra- and inter-series dependencies with causality and scalability maintained. In Table 1, we list representative time-series Transformers and highlight their differences.
Table 1: Comparison among representative time-series Transformers.
Model PatchTST iTrans. TimeXer UniTST Moirai Timer Timer-XL
(2022) (2023) (2024b) (2024a) (2024) (2024c) (Ours)
Intra-Series
Inter-Series
Causal Trm.
Pre-Trained
@@ 3 Approach
In this section, we first introduce a decoder-only Transformer to illustrate the procedure of next token prediction on univariate time series. As an extension, we design TimeAttention and propose Timer-XL for unified time series forecasting. It is applicable to univariate, multivariate, and covariate-informed scenarios by generalizing the context from 1D sequences to 2D time series.
@@ 3.1 Timer
Timer (Liu et al., 2024c) is a time-series Transformer trained by next token prediction (Bengio et al., 2000), which regards single-dimensional time series as non-overlapping patch tokens.
@@
For simplicity, we omit the block index $l$. Timer adopts $\mathbf{W}_{e},\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ that independently embed and project the token embeddings as $\mathbf{H}=\{\mathbf{h}_{i}\}\in\mathbb{R}^{T\times D}$. $\operatorname{TrmBlock}$ includes feed-forward network and self-attention with the temporal causal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$. $\mathbf{h}_{i}\in\mathbb{R}^{D}$ is the context representation of the previous $i$ tokens. All predicted $\hat{\mathbf{x}}_{i+1}$ are supervised with ground truth via MSE loss.
@@ 3.2 Generalize 1D Sequences to 2D Time Series
Compared with Equation 2, the multivariate context length increases from $T$ to $NT$. By contrast, the benefit is that this paradigm learns causal dependencies within each sequence while incorporating exogenous variable correlations from other sequences, making it a universal forecasting paradigm that outperforms channel-independent (Nie et al., 2022) or variable-centric models (Liu et al., 2023).
Technically, we independently apply $\mathbf{W}_{e}\in\mathbb{R}^{D\times P}$ on each token to obtain patch-wise representation $\mathbf{h}_{m,i}\in\mathbb{R}^{D}$, which will encompass contextual information from $Ni$ tokens through Transformer blocks and be eventually projected by $\mathbf{W}_{d}\in\mathbb{R}^{D\times P}$ into the predicted patch token $\hat{\mathbf{x}}_{m,i+1}$.
@@
Position embedding has not been sufficiently explored in time-series Transformers. To avoid inherent permutation-invariance of self-attention, positional embedding is required to reflect the chronological order of tokens on the temporal dimension. As for the variable dimension, shuffling the input order of variables should not affect anything other than the output order of variables. Formally, the processing on multiple variables should be permutation-equivalent (Zaheer et al., 2017).
@@
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times d_{k}}$ and $d_{k}$ is the dimension of the query, key, and value. $\mathbf{R}_{\theta,t}\in\mathbb{R}^{d_{k}\times d_{k}}$ is the rotary matrix with rotation degree $t\cdot\theta$, $\mathds{1}{(}\cdot)$ is the indicator function, and $u,v\in\mathbb{R}$ are learnable parameters for the token to distinguish its endogenous and exogenous time series.
@@
We provide an intuitive example to illustrate the causal dependencies within multivariate time series: considering the 2nd token of time series A. To predict its next token, its representation $\mathbf{h}$ should be exactly dependent on the tokens-$\{1,2,4,5\}$. Similarly, we provide all causal dependencies of each token in Figure 12. Based on the visualized attention mask and variable dependencies presented in Figure 2, where all variables are inter-correlated, all token dependencies in $\mathcal{A}$ can be formally disentangled by the Kronecker product into (1) the adjacency matrix of the variable dependency graph $\mathcal{C}\in\mathbb{R}^{N\times N}$ and (2) the causal temporal mask $\mathcal{T}\in\mathbb{R}^{T\times T}$:
@@
Eventually, token representations in $\mathbf{H}=\{\mathbf{h}_{m,i}\}\in\mathbb{R}^{NT\times D}$ will be independently processed by feed-forward network and layer normalization, and fed into the next Transformer block.
@@
In multivariate forecasting, the variable dependency forms the complete graph, presenting an all-one matrix $\mathcal{C}$. By generalizing TimeAttention on multiple sequences, Transformers can leverage its length-flexibility to encompass relevant covariates as well. In this case, Timer-XL is adapted in two steps: (1) formulate the customized variable dependency as $\mathcal{C}$ and (2) optimize the model using the supervision of target variables. An example (target-$A$-covariate-$B$) of TimeAttention is illustrated on the right of Figure 2. In a nutshell, we adopt position embeddings for the temporal and variable dimensions. To achieve unified time series forecasting, we flatten 2D time series into a unified context and capture fine-grained causal token dependencies.
@@ 4 Experiments
We conduct evaluations of Timer-XL in three aspects, including (1) supervised training as a task-specific forecaster, (2) large-scale pre-training as a zero-shot forecaster, and (3) assessing the effectiveness of TimeAttention and model efficiency. Given that the long-context forecasting paradigm receives less attention in the community, which can be concealed due to the performance saturation on previous benchmarks (Makridakis et al., 2020; Wu et al., 2022), we established new long-context forecasting benchmarks. Detailed experimental configurations are provided in Appendix B.